OpenStudy (anonymous):

The sum of the height and the radius of the base of a right circular cone is 9 inches. What is the maximum volume of the cone?

OpenStudy (anonymous):

So first, we have the equation for the volume of a cone, \[V = \frac{ \pi r^2 h }{ 3 }\] Then, you were given an additional constraint that, \[h+r = 9\] Thus, we know that, \[h=9-r\] Can you work it out from there?